# GroundGazer: Camera-based indoor localization of mobile robots with millimeter accuracy at low cost

### v4

@@ Abstract
Highly accurate indoor localization systems with mm positioning accuracy are currently very expensive. They include laser trackers, total stations, and motion capture systems relying on multiple high-end cameras. In this work, we introduce a high-accuracy, planar indoor localization system named GroundGazer (GG) for autonomous mobile robots (AMRs). GG estimates the AMR’s position with mm and its heading with sub-degree accuracy. The system requires only a monocular (fisheye) camera, a chessboard floor, and an optional laser diode. Our system is simple and low-cost, easy to set up, portable, robust, scalable to large areas and robot swarms, and extendable to 3D position and orientation estimation.
@@ I Introduction
Benchmarking lower-accuracy indoor positioning systems with up to cm localization accuracy requires localization approaches with reliable mm positioning accuracy. Currently, such systems are very costly (generally $1000\,\mathdollar$, with costs rising quickly with increasing size of the coverage area). In this work, we introduce a novel camera-based indoor localization system for AMRs which achieves mm localization and sub-degree heading accuracy at multitudes lower cost than existing systems of similar precision. We name it GroundGazer, as the camera placed on the AMR gazes at the passive position references on the ground.
Before describing the GG system, we give an introduction into currently existing and researched indoor localization technologies for AMRs with mm positioning accuracy. For an overview of general indoor localization technologies, we refer to the surveys in [24] and [34]. Comprehensive overviews of optical indoor localization systems can be found in [22] and [25].
I-A Localization technologies with mm positioning accuracy
Motion capture systems consist of passive infrared (IR) markers placed on the to be tracked AMR. With one marker, 3D positioning is possible. A minimum of three non-collinear markers enable 3D position and orientation estimation. Motion capture systems like the commercial systems by Qualisys and Vicon [^1] Available at: https://www.qualisys.com/ and https://www.vicon.com/. are very popular in robotics and industry. While capable of sub-mm 3D localization and sub-degree orientation estimation, they require line of sight (LOS), multiple high-end IR cameras, and become very expensive if large areas are to be observed. Further, the IR cameras not installed around the AMR complicate the time synchronization when the motion capture system is used as reference for another localization system with sensors installed on the AMR. Highly accurate time synchronization is required, since otherwise the mm accuracy is lost through badly matched time stamps.
Total stations (tachymeters) pose a another solution for 3D position estimation with mm or sub-mm accuracy, e.g. used as 3D position reference system in [11]. Total stations enable long range sensing, but like motion capture systems, they are expensive and require LOS and more complex time synchronization mechanisms.
A novel sub-mm positioning system with long distance capabilities based on radio signals and interferometry that has recently seen real-world deployments is developed by Koherent [^2] Available at: https://koherent.io/.
Radio frequency (RF)-based backscatter systems using frequency modulated continuous wave (FMCW) radar and (semi-)active reflective tags (e.g. Van-Atta arrays) that modulate the reflected radar signal can also achieve 3D mm localization accuracy using high-precision range estimates of the tags and multilateration (MLAT) [1] or triangulation [33] for positioning. The system in [1] could also demonstrate good performance with large radar-tag distance.
In acoustic ultrasound systems, small timing errors in the measured time of flight (ToF) of the signals or in the synchronization between anchors cause lower positioning errors than in RF-based ultra wideband (UWB) systems. The reason lies in the much lower propagation speed of acoustic waves. The company ZeroKey claims real-time and up to 1.5 mm 3D positioning accuracy with their commercial Quantum RLTS 2.0 system, given 6 acoustic anchors with ideal geometric placement [35]. Positioning is performed with multiple active acoustic anchors and ToF measurements between anchors and the to be tracked AMR and by applying MLAT for localization.
While ToF-based localization with UWB systems only gives dm to cm accuracy, recent studies show mm accuracy by combining initial, coarse ToF-based range estimates with refined, ambiguity-resolved phased-based range estimates [21].
Deep learning-based channel state information (CSI) fingerprinting with a massive multiple input multiple output (MIMO) system could recently achieve mm accuracy in a small area and under LOS conditions [20]. However, the performance of such system degrades with changes in the environment, causing domain shifts between the measured fingerprints during inference and the pre-recorded fingerprint database.
Fiducial markers pose another solution for highly precise localization. In [14], the combination of high-end binocular cameras and custom markers achieves sub-mm localization accuracy. Downsides are again the high cost of the cameras and the required LOS. The authors of [4] proposed a docking system with sub-mm 2D positioning accuracy using quick response (QR) codes and an RGB camera. Although low-cost, the very small camera-QR code distance and the need for many QR codes limit its use for general AMR localization.
Other systems with (sub-)$m$ accuracy based on visual markers, laser projections, magnetic fields, laser trackers, pseudolites, and infrared laser beams (iGPS) can be found in [22] and [23].
I-B Contributions
@@
We introduce a novel mm accuracy localization approach with the intended application to serve as reference for less accurate indoor positioning systems. The GG prototype described in this paper requires only a chessboard floor (or similar grid pattern), a camera (here a global shutter RGB fisheye camera), and an optional laser diode, i.e. low-cost off-the-shelve components. Currently, it only allows for planar position ($x$, $y$) and heading $\theta$ estimation, which is sufficient for our application as reference system for planar AMR position and heading estimation with known or irrelevant height, roll and pitch. In Sec. VI, we discuss an adaption that yields 3D position and orientation estimates, but has higher computational complexity and possibly lower accuracy.
@@
The idea for our GG system originates from [3]. They developed a fiducial marker localization system that requires a chessboard floor in addition to the fiducial markers, where a car is equipped with multiple fiducial markers placed on a plate mounted at the front of the car. To evaluate the positioning accuracy of this system, they installed two point laser diodes that project two points from the front and the back of the car onto the chessboard floor beneath it. The corner coordinates of the chessboard serve as position references and they measure the two laser point locations on the floor as ground truth by hand. The laser point at the front of the car is used as vehicle reference position and the heading is derived from the orientation of the line connecting both points w.r.t. to the chessboard. Our GG system builds on that idea, resulting in a highly accurate, camera-based indoor localization system.
I-C System overview
Our localization system is depicted in Fig. 1(a).
@@ I-C System overview
((a)) Our system in the laboratory equipped with a chessboard floor and QR codes. The AMR has a ground facing camera and a laser diode that projects a crosshair on the floor at the front.
@@
It consists of an AMR, which is equipped with an RGB fisheye camera ($100\,\mathdollar$) with large field-of-view (FoV) and global shutter. Global shutter cameras have lower motion blur than rolling shutter cameras due to parallel collection of all pixels. The camera is facing the ground and a green laser diode (negligible cost) projects a crosshair on said ground. The ground itself is made up of a chessboard pattern with well known grid size.
@@
The localization system works by detecting the crosshair in a square of the chessboard. We specifically selected a blue and white chessboard floor and a green laser diode, as green lasers are most easily visible and other commercially available black and white chessboard floors complicate the crosshair detection, caused by the large difference in laser reflective behavior between black and white squares. The camera image is calibrated such that the heading of the AMR can be estimated based on the orientation of the chessboard in the image. The corner coordinates of the chessboard squares are known and serve as passive position references for AMR localization.
@@
Since the chessboard corner coordinates are ambiguous to the camera which only observes a small, local area of the periodic chessboard pattern, we additionally equip the floor with different QR codes as global position and orientation references. All QR codes are oriented in the same direction, parallel to the chessboard lines, and their corners can therefore be used to extract the $x$- and $y$-directions of the chessboard lines in global coordinates. The QR code IDs allow for identification of the QR code squares and global positioning. By measuring the distance in $x$- and $y$-direction between a detected QR code square center and the center of the square in which the crosshair lies (neglecting the trivial case where they are identical), we can further identify the crosshair square and its known corner coordinates without requiring a QR code in each square. Measuring the relative position of the crosshair w.r.t. its square corners (at least three corners are required to build a coordinate system, but we use four corners as known from a detected square) gives the AMR’s position and the angle between the chessboard lines in $x$-direction and the vertical image direction gives its heading. The AMR position is assumed as the crosshair position. An image describing the AMR positioning is shown in Fig. 1(b).
High position and heading accuracy is enabled by placing the camera close to the ground and by precise estimation of the chessboard square corners, the chessboard lines, and the crosshair center. For this prototype, a camera with fisheye lens was chosen because of their large FoV, allowing observation of multiple chessboard squares with a small camera-ground distance. The whole processing chain, which transforms the initial fisheye image, detects the QR codes, the crosshair, and the chessboard square corners, and finally estimates the AMR’s position and heading, is described in Sec. III.
II Related works with ground gazing cameras
@@
In this section, we focus on closely related systems that employ ground gazing cameras for global localization. A general overview of indoor positioning systems with (sub)-mm accuracy for AMRs was given in Sec. I-A.
One approach to global localization using ground gazing cameras is by avoiding external infrastructure, i.e. localization using only the given ground texture. Relative positioning can be realized through simultaneous localization and mapping (SLAM) [32], but we are interested in absolute, global positioning. Such positioning is possible with feature-matching of images collected during inference with a pre-collected image database of the ground [17, 36, 7, 12, 6]. The advantage of this is the applicability in (ideally) any kind of environment without requiring new infrastructure, making it a practical method e.g. for warehousing. However, feature-matching requires an expensive measurement campaign for collection of reference images of the floor with accurate ground truth, e.g. from motion capture systems. These approaches have also shown varying reliability for different surfaces [36] and the position errors are in the low cm range (ignoring outliers, but also not yet using temporal correlations) and the heading errors are in the low degree range [6]. Positioning errors further seem to increase with coverage area (map size) [6]. The commercial feature-matching-based systems by Accerion are reported to deliver sub-mm accuracy, but only for repeated tasks such as docking, where the task was accurately executed before, while stating cm accuracy for global localization [^3] Due to the limited product specifications on the company website (no published data sheets), we are relying on the information given on their Youtube channel found at https://www.youtube.com/@accerion4564.. For relative localization, features of successive images can similarly be matched to build a vision-based odometer [27]. While feature-matching approaches require no additional infrastructure, we argue that using known position references is the more suitable approach for reference systems due to higher accuracy, robustness, easier setup, and lower hardware requirements.
Various approaches have considered placing infrastructure on the floor for absolute, global localization. In [37], ArUco markers are distributed on the floor and combined with odometry in an Extended Kalman filter (EKF) for AMR localization. However, they don’t give localization accuracy results. AMR localization with a simulated grid of ApriTags yields cm accuracy in [18]. QR codes are also commonly applied, e.g. for a heeled robotic walking assistant in [26], achieving dm accuracy with additional odometry and gyroscope data. In [16], a drone with a ground facing camera is localized with a QR code grid on the floor and using the nearest QR code for position and heading estimation, showing cm accuracy. Different fiducial marker types are benchmarked regarding their detection behavior in ground-based localization in [8].
It should be noted, however, that we believe low-cost mm accuracy with floor installations of ArUco, AprilTag, QR codes, or similar fiducial markers is possible. With ceiling installations of ArUco markers and much larger camera-marker distances, we could achieve low cm accuracy in previous work [15] and [4] reports sub-mm accuracy with very small camera-QR code distance. Similar commercial systems can also be found. One example is the company Hikrobot, which provides AMRs and vision-based localization solutions using a visual marker grid on the floor for warehousing, logistics etc. [^4] Available at: https://www.hikrobotics.com/en/. The Hercules AMR by Amazon robotics is another commercial system using such floor marker grid for navigation.
Opposed to existing systems, we can deliver absolute 2D position and heading estimates with mm and sub-degree accuracy at low cost and with minimal setup. Our processing can likely also be sped-up through further optimization as discussed in Sec. VI to enable near real-time mm accuracy during online processing with common AMR velocities of $1\text{\,}\frac{m}{s}$, factoring in additional delays from exposure time and image transmission to the processor ($10\text{\,}\mathrm{m}\mathrm{s}$ delay at $1\text{\,}\frac{m}{s}$ causes $10\text{\,}\mathrm{m}\mathrm{m}$ error, i.e. cm online accuracy). This online bottleneck might be solved by integrating a high-frequency inertial measurement unit (IMU). Neither current SLAM nor feature-matching approaches can produce this due to higher processing times. To the best of our knowledge, no system like GG exists right now for indoor localization of AMRs that offers a similar combination of high accuracy, low cost, simplicity, easy setup, scalability, portability, and fast processing.
@@ III-A Image transformation
The first step of our processing is the transformation of the distorted and skewed image collected by the fisheye camera to a top-down view of the chessboard floor. This step consists of two transformations. First, the fisheye image needs to be undistorted. We apply the built-in undistort function from OpenCV [5] [^5] If not specified otherwise, we always refer to OpenCV functions for the remainder of the paper.. This function takes as input the distorted fisheye image, the intrinsic matrix $\mathbf{K}$ of the camera, and its distortion coefficient vector $\underline{d}$, and outputs the undistorted image. The intrinsic matrix $\mathbf{K}$ and the distortion coefficient vector $\underline{d}$ have to be determined beforehand by camera calibration. We used a printed chessboard pattern and the calibrateCamera function for this task. A distorted fisheye image and the result after undistortion are shown in Fig. 2(a) and Fig. 2(b), respectively. The second step transforms the undistorted image to a top-down view. This can be realized with the warpPerspective function, which takes the undistorted image and the homography matrix as input and gives the top-down view. The homography relates the undistorted and the top-down view. It can be estimated offline with the findHomography function and corresponding image points of the two views. We define a square of size (300, 300) pixels with horizontal and vertical edges for the wanted top-down view after perspective transformation and we select the four corresponding chessboard corner points from the undistorted image. The perspective transformed top-down image is given in Fig. 2(c). The AMR orientation has to match the direction of the calibration square, i.e. in our case $\theta=90^{\circ}$, to align the heading with the transformed chessboard orientation used for heading estimation. Alternatively, one could use the angle between a similarly calibrated, vertical crosshair line that points in heading direction and the $x$-direction chessboard line. This option would make heading estimation independent of the camera orientation and thus of camera vibrations. However, it is still susceptible to lesser laser diode vibrations and it makes heading estimation dependent on the crosshair line detection, which is worse than the chessboard line detection due to broader edges, crosshair distortion on uneven ground etc., and therefore not chosen.
@@
Figure 2: Transformation of an original image recorded with an AMR heading slightly below $90^{\circ}$ in a) through undistortion in b) and perspective transformation in c).
@@ III-B QR code detection and masking
QR codes are detected using OpenCV, which outputs the four corner points of the QR code in known order. This gives the global orientation information, previously shown in Fig. 1(b). The QR code ID is extracted with the pyzbar library [^6] Availabe at: https://pypi.org/project/pyzbar/., as it gave better results than OpenCV. The QR ID enables identification of the respective chessboard square. By computing the difference in $x$ and $y$ between the square centers where the QR code is detected and where the crosshair is detected, the QR ID also gives us the square in which the crosshair lies and thus global localization capabilities.
Since the QR codes would disturb the successive image processing operations, we mask them out. Therefore, the average color around the detected QR code is computed and the region of the QR code is replaced with a slightly larger rectangle of this color. This masking procedure might also mask out parts of the crosshair if it lies on a QR code. The QR code masking could be improved in future work, but our experiments have shown that our system can cope with irregular missed crosshair detections. By proper parametrization of the Hough transform [10] in the crosshair detection, we can also detect an incomplete crosshair.
III-C Crosshair detection and masking
The green crosshair can be extracted by filtering the image with a mask that keeps the green image parts. This mask has to be carefully tuned, as both the different colors of the chessboard and varying illumination condition influence the detected crosshair color. We therefore apply a concatenation of multiple masks. The first mask filters out all pixels with low green values. We then convert the image to the HSV color space, which was more robust than only working in the RBG space. Three additional masks filter out pixels with low saturation, low hue, and pixels where the sum of hue and value is low. A combined crosshair mask is found in Fig. 3(a). The horizontal and vertical lines in a region of 150 pixels around the estimated crosshair position from the previous time step are then found by Hough transformation, clustered into lines with similar angle and start- and endpoints, and the start and end $x$-values of the clustered lines are averaged for the vertical, and the $y$-values for the horizontal line. The crosshair location in image coordinates is then estimated as the intersection of the horizontal and vertical crosshair lines. The final crosshair lines and the detected crosshair position are shown in Fig. 3(b). To not disturb the chessboard square detection, we mask out the crosshair with an extended and broadened version of the detected crosshair.

### v5

# GroundGazer: An optical reference system for planar localization with millimeter accuracy at low cost

@@ Abstract
Highly accurate indoor localization systems with absolute mm positioning accuracy are currently expensive. They include laser trackers, total stations, and motion capture systems relying on multiple high-end cameras. In this work, we introduce a high-accuracy, planar indoor localization system named GroundGazer (GG) for autonomous mobile robots (AMRs). GG estimates the AMR’s planar position with mm and its heading with sub-degree accuracy. The system requires only a monocular (fisheye) camera, a chessboard floor, and an optional laser diode. Our system is simple and low-cost due to the chessboard floor, robust, scalable to multiple robots, and extendable to 3D position and orientation estimation.
@@ I Introduction
Benchmarking lower-accuracy indoor positioning systems with up to cm localization accuracy requires localization approaches with reliable mm accuracy. Currently, such systems (described in Sec. II-B) are very expensive. Costs for deploying a motion capture system, a common solution for accurate position referencing, can easily exceed $\mathdollar 10000$, even for a small laboratory. While some alternatives might be cheaper, they usually come with other drawbacks such as strict line of sight (LOS) requirements, or degradation with multi-path.
I-A Contributions
@@
In this work, we introduce a camera-based, planar indoor localization system for AMRs which achieves mm localization and sub-degree heading accuracy at multitudes lower cost than existing systems of similar precision for small rooms. We name it GroundGazer, as the camera placed on the AMR gazes at the passive position references on the ground. Its intended application is to serve as reference for less accurate localization methods. The proposed GG prototype requires only a chessboard floor (or similar grid pattern), a camera (here a global shutter RGB fisheye camera), and an optional laser diode, i.e. low-cost, off-the-shelve components. The current implementation estimates planar position ($x$, $y$) and heading $\theta$, which is sufficient for planar AMR localization with known or irrelevant height, roll, and pitch. In Sec. VI, we discuss an extension to 3D position and orientation estimation, at the cost of higher computational complexity and possibly reduced accuracy.
@@
The idea for GG originates from [2]. To evaluate the accuracy of their proposed positioning system, they installed two point laser diodes that project two points from the front and the back of their localized vehicle onto the chessboard floor beneath it. The manually collected positions of those two points w.r.t. the known corner coordinates of the chessboard enable precise position and heading referencing. Our GG system relies on the same idea, resulting in a highly accurate, automated, camera-based indoor localization system.
I-B System overview
The GG prototype is depicted in Fig. 1(a).
@@ I-B System overview
((a)) Our system in the laboratory equipped with a chessboard floor and QR codes. The AMR has a ground gazing camera and a laser diode that projects a crosshair on the floor at the front.
@@
[…]
@@
The localization system works by detecting the crosshair in a chessboard square. We selected a blue and white chessboard floor and a green laser diode, as green lasers are most easily visible and other commercially available black and white chessboard floors complicate the crosshair detection, caused by the large difference in laser reflective behavior between black and white squares. The camera image is calibrated such that the orientation of the chessboard in the image directly gives the heading. The corner coordinates of the chessboard squares are known and serve as passive position references for AMR localization.
@@
Since the camera only observes a small, local area of the periodic chessboard pattern, the corner coordinates are ambiguous. Adding different QR codes as global position and orientation references to the floor resolves this ambiguity. All QR codes are coarsely aligned with the chessboard lines, and their corners can thus be used to extract the $x$- and $y$-directions of the chessboard lines in global coordinates. The QR code IDs allow for identification of the QR code squares and global positioning. By measuring the offset between a detected QR code square and the square containing the crosshair, the crosshair square and its known corner coordinates can be identified without requiring a QR code in each square.
The AMR position is defined as the crosshair center location P5. Its planar position $(x,y)$ is given by the relative position of the crosshair w.r.t. its square corners P1-P4 and the heading $\theta$ is given by the angle between the chessboard lines in $x$-direction and the vertical image axis. An image describing the AMR positioning is shown in Fig. 1(b).
GG’s system design enables high accuracy with simple visual infrastructure. While the requirement of a chessboard floor limits GG’s use as a general localization method, deploying such chessboard floor comes at reasonable effort and at low cost for small areas. GG is not restricted to visible light and could also be realized with an active infrared (IR) camera and a passive IR grid, improving robustness to ambient illumination and allowing the positioning infrastructure to remain largely unobtrusive for human users. The image processing to estimate the AMR’s position and heading is described in Sec. III.
II Related works
@@
We focus on closely related systems in this section. For an overview of general indoor localization technologies, we refer to the surveys in [14] and [21], and for optical indoor localization systems to [12] and [15].
II-A Related works with ground gazing cameras
Infrastructure-free approaches localize using only the given ground texture. Relative positioning can be realized through simultaneous localization and mapping (SLAM) or feature-matching in successive images to build a vision-based odometer [17]. Absolute positioning is possible with feature-matching of images collected during inference with a pre-collected image database of the ground [23, 4, 20]. However, mapping a floor and anchoring it to the real world with correct absolute scale requires an expensive measurement campaign with accurate ground truth, e.g. from motion capture systems [4].
While feature-matching requires no infrastructure, using known position references is the more suitable approach for reference systems due to higher accuracy, robustness, easier setup, and lower hardware requirements. Various systems place infrastructure (fiducial markers) on the floor for absolute, global localization of AMRs [9], and other robots, such as walking assistants [16] and drones [8]. By applying popular markers including QR codes, ArUco markers, or AprilTags, they achieve dm to cm localization accuracy. A QR code-based docking system in [3] reports sub-mm accuracy with very small camera-QR code distance. Although low-cost, the very small camera-QR code distance and the need for many QR codes limit its use for AMR localization. Visual floor marker grids for AMR navigation can also be found in commercial systems, e.g. the Hercules AMR by Amazon robotics.
The chessboard line grid in GG is easier to realize than a precisely placed marker grid and offers more accurate corner detection with high robustness against occlusion due to its redundant pattern. In a laboratory, it can also facilitate installation of other equipment.
II-B Localization technologies with mm positioning accuracy
Optical systems include motion capture systems like the commercial systems by Qualisys and Vicon [^1] Available at: https://www.qualisys.com/ and https://www.vicon.com/.. They consist of passive IR markers placed on the tracked AMR. With one marker, 3D positioning is possible. A minimum of three non-collinear markers enable 3D position and orientation estimation. While capable of sub-mm 3D localization and sub-degree orientation estimation, they require LOS, multiple high-end IR cameras, and become very expensive for large-scale deployments. Further, since the cameras are not installed around the AMR, accurate time synchronization is required to ensure mm accuracy. A different optical solution are fiducial markers. In [7], sub-mm accuracy is achieved, but only with costly cameras and metrology-grade equipment to precisely install custom reference markers. The previously mentioned QR code docking system in [3] also gives sub-mm accuracy, but is limited by camera-QR code distance.
Total stations (tachymeters) pose another solution for 3D position estimation with mm or sub-mm accuracy, e.g. as used in [5]. Total stations enable long range sensing, but like motion capture systems, they are expensive and require LOS and accurate synchronization.
In acoustic ultrasound systems, the low propagation speed of sound greatly reduces sensitivity to timing and synchronization errors compared to radio frequency (RF)-based ultra wideband (UWB) systems. Using time of flight (ToF)-based measurements between multiple active anchors and the tracked user, a commercial system reports real-time 3D positioning accuracy of up to 1.5 mm using 6 acoustic anchors with ideal geometric placement [22].
While ToF-based localization with UWB systems only gives dm to cm accuracy, recent studies show mm accuracy by combining initial, coarse ToF-based range estimates with refined, ambiguity-resolved phased-based range estimates [11]. Other RF-based systems include the interferometry-based system developed by Koherent [^2] Available at: https://koherent.io/, and frequency-modulated continuous wave (FMCW) radar-based backscatter systems using reflective tags that modulate the reflected signal [1, 18]. Deep learning-based channel state information (CSI) fingerprinting with a massive multiple input multiple output (MIMO) system could recently also achieve mm accuracy in a small area and under LOS conditions [10]. However, domain shifts between training and inference, caused by environmental changes, reduce the accuracy of such systems. RF-based systems often require specialized hardware and degrade under multipath. At smaller wavelength, using light detection and ranging (LiDAR) with reflectors at known positions as references has also shown mm accuracy [19], but LiDAR has high sensor cost.
Additional systems with (sub-)$m$ accuracy based on visual markers, laser projections, magnetic fields, laser trackers, pseudolites, and infrared laser beams (iGPS) can be found in [12] and [13].
@@ III-A Image transformation
The initial step of our GG localization algorithm transforms the distorted and skewed image collected by the fisheye camera to a top-down view of the chessboard floor. This step consists of two transformations. First, the fisheye image is undistorted. The required intrinsic matrix $\mathbf{K}$ of the camera and its distortion coefficient vector $\underline{d}$ are estimated with a printed chessboard pattern. A distorted fisheye image and the undistorted result are shown in Fig. 2(a) and Fig. 2(b), respectively. Then, the undistorted image is transformed to a top-down view by perspective transformation, depicted in Fig. 2(c). Since GG localizes the crosshair center relative to its square corners, errors in undistortion and perspective transformation have little influence on the positioning accuracy. The homography matrix for perspective transformation is estimated offline with four corresponding square corner points of both views. The AMR heading has to match the orientation of the transformed chessboard to align them. We calibrate at $\theta=90^{\circ}$, thus the edges of the top-down view square are parallel to the image axes. We set its size to (300, 300) pixels. The four corresponding points from the undistorted image view are the corners of the central crosshair square.
Alternatively, heading could be estimated with the angle between a similarly calibrated, vertical crosshair line that points in heading direction and the $x$-direction chessboard line. This makes heading estimation independent of the camera orientation and thus of camera vibrations. However, it is still susceptible to lesser laser diode vibrations and it makes heading estimation dependent on the crosshair line detection, which is worse than the chessboard line detection due to broader edges, crosshair distortion on uneven ground etc., and therefore not chosen.
@@
Figure 2: Transformation of an original image in a) through undistortion in b) and perspective transformation in c).
@@ III-B QR code detection and masking
QR code detection gives the four corner points of a QR code in known order and its ID, which enables identification of all chessboard squares and their corners.
As the QR codes would disturb successive image processing, we mask them out. Our procedure might also mask parts of the crosshair if it lies on a QR code and could be improved in future work. Nevertheless, our experiments have shown that GG can cope with irregular missed crosshair detections and by proper parametrization of the Hough transform in the crosshair detection, we can also detect an incomplete crosshair.
III-C Crosshair detection
The green crosshair can be extracted by filtering the image with a mask that keeps the green image parts. This mask has to be carefully tuned, as both the different colors of the chessboard and varying illumination condition influence the detected crosshair color. We therefore apply a concatenation of multiple masks. The first mask filters out all pixels with low green values. We then convert the image to the HSV color space, which was more robust than only working in the RBG space. Three additional masks filter out pixels with low saturation, low hue, and pixels where the sum of hue and value is low. A combined crosshair mask is found in Fig. 3(a). The horizontal and vertical lines in a region of 150 pixels around the estimated crosshair position from the previous time step are then found by probabilistic Hough transformation, clustered into lines with similar angle and start- and endpoints, and the start and end $x$-values of the clustered lines are averaged for the vertical, and the $y$-values for the horizontal line. The crosshair location in image coordinates is then estimated as the intersection of the horizontal and vertical crosshair lines. The crosshair lines and the detected crosshair position are shown in Fig. 3(b).